OpenStudy (anonymous):

differentiate. 2^sin(pi)x .... im completely lost. please explain steps.

OpenStudy (anonymous):

does this really say \[2^{\sin(\pi)x}\]?

OpenStudy (anonymous):

\[2^{\sin(\pi x)}\] perpaps

OpenStudy (anonymous):

or is it \[2^{\sin(\pi x)}\]

OpenStudy (anonymous):

flip the pi and x

OpenStudy (anonymous):

if it is the second one, recall that \[\frac{d}{dx}2^x=2^x\ln(2)\] so by the chain rule \[\frac{d}{dx}2^{\sin(\pi x)}=2^{\sin(\pi x)}\times \ln(2)\times \cos(\pi x)\times \pi\]

OpenStudy (anonymous):

oh wait . you got it right. my basd vison

OpenStudy (anonymous):

ok got it. thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hello myininaya!

myininaya (myininaya):

satellite i need your help

OpenStudy (anonymous):

i see you in chat

myininaya (myininaya):

ok